OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope = ; y-intercept = (0, 2) slope = -; y-intercept = (0, 2) slope = ; y-intercept = (0, -6)

OpenStudy (anonymous):

3y = 5x + 6 y = 5/3 x + 2 slope = 5/3 and y intercept = (0,2)